1. Suppose that TapDance, Inc.’s, capital structure features 65 percent equity, 35 percent debt, and that its before-tax cost of debt is 7 percent, while its cost of equity is 12 percent. Assume the appropriate weighted average tax rate is 34 percent.

What will be TapDance’s WACC? (Round your answer to 2 decimal places.)

2. Suppose that MNINK Industries’ capital structure features 63 percent equity, 7 percent preferred stock, and 30 percent debt. Assume the before-tax component costs of equity, preferred stock, and debt are 11.80 percent, 9.70 percent, and 9.00 percent, respectively.

What is MNINK’s WACC if the firm faces an average tax rate of 34 percent? (Round your answer to 2 decimal places.)
Answer #1

1.After-tax cost of debt=7*(1-tax rate)

=7*(1-0.34)=4.62%

WACC=Respective costs*Respective weight

=(4.62*0.35)+(0.65*12)

=9.42%(Approx).

2.After-tax cost of debt=9*(1-tax rate)

=9*(1-0.34)=5.94%

WACC=(5.94*0.3)+(0.63*11.8)+(0.07*9.7)

=9.90%(Approx).